OpenStudy (anonymous):

A cube has 6 faces, 8 vertices, and 13 edges. True False 2

OpenStudy (phi):

can't you find a box and count them? But I am suspicious of the 13 edges.

OpenStudy (anonymous):

SAME HERE! the 13 edges is what confused me

OpenStudy (anonymous):

you can google how many edges it should have...

OpenStudy (anonymous):

12 edges

OpenStudy (anonymous):

Faces + Vertices= Edges.

OpenStudy (phi):

Faces + Vertices= Edges+2

OpenStudy (anonymous):

so, it would be 6+8=14+2. from that point i am confused. is it 14 edges?

OpenStudy (anonymous):

or 16 edges?

OpenStudy (phi):

There are 6 faces: top and bottom and 4 sides. Look at a box. there are 8 vertices (corners): 4 at the top and 4 at the bottom = 8 there are 12 edges: 4 at the top, 4 at the bottom, and 4 for the sides 6+8= 12+2

OpenStudy (anonymous):

got it! @phi thank you for the clarification